Reece:

When I looked outside, I saw birds and cats. I counted 19 heads and 42 legs. ( I counted because I thought: This would be a good math problem!) How many birds and how many cats did I see in the yard

Reece:

help plz

Reece:

hey where did ur work go

Reece:

why???

Reece:

do u know how to do it

freshavacado850:

yes i just did somthing wrong

Reece:

ok plz help

Angle:

reece, are you able to do algebra yet? sometimes problems like these are asked before people learn algebra, so I'd like to ask

Reece:

no

Reece:

i am confused on how to do this problem

Angle:

ok, no problem here is the strategy though The (number of birds) + (number of cats) = 19 heads also, because birds have two legs and cats have four legs each (2 legs per bird)x(number of birds) + (4 legs per cat)x(number of cats) = 42 legs total

Angle:

I know I just typed a lot, so read through it and let me know if you don't understand

Reece:

idon't understand how to do it

Angle:

b + c = 19 animals 2b + 4c = 42 legs these are our two equations

Reece:

ok I figure it out I think

Reece:

but can you make sure its right plz

Angle:

of course!

Reece:

theres 7 birds (14 legs) and 9 cats (36 legs)

Reece:

is this right???

Angle:

hmmm... I don't think 7 birds and 9 cats adds up to 19 animals?

Reece:

ok give me another sec. lol

Reece:

ok so 7 cats and 11 birds?????

Reece:

right??

Angle:

11+7=18 but you're getting closer

Reece:

ok 13 birds and 6 cats right???

Angle:

13 birds = 26 legs 6 cats = 24 legs 26 + 24 = 50

Angle:

let's try even more birds

Reece:

omg !!!!! ok give me another sec

Reece:

ok 3 cats and 15 birds????

Angle:

oof so so sooo close! 3 + 15 = 18 not 19 though

Angle:

just a couple more birds I think

Reece:

17 birds and 2 cats right???

Angle:

perfect !!! you got it :) :)

Reece:

omg thank you sooooooo much

freshavacado850:

good jobbb

Reece:

thx

Angle:

Thanks for not giving up!

Angle:

I know that took a lot of effort

Reece:

lol thats thx to you and thank you for not giving up on me